OpenStudy (anonymous):

Verify the trig identity: cot t + sin t/ 1 +cos t = csc t

OpenStudy (anonymous):

\[\cot (t)+\sin(\frac{ t }{ 1 })+\cos(t)= \csc(t)\]

OpenStudy (anonymous):

\[\csc (t)=\frac{ 1 }{ \sin(t) }\] \[\cot (t)= \frac{ \cos(t) }{ \sin(t) }\]

OpenStudy (anonymous):

Sorry, I still don't get it.

OpenStudy (anonymous):

\[\frac{ \cos(t) }{ \sin (t) }+\sin(t) +\cos (t)= \frac{ 1 }{ \sin(t) }\]

OpenStudy (anonymous):

common denominator is sin(t)

OpenStudy (anonymous):

\[\frac{ \cos(t) }{ \sin(t) }+\frac{ \sin(t)*\sin(t) }{ \sin(t) }+\frac{ \cos(t)*\sin(t) }{ \sin(t) }=\]

OpenStudy (anonymous):

something wring with the equation ?!

OpenStudy (anonymous):

i couldn't come up with the answer, so i evaluate both right and left sides they are not not equal

OpenStudy (anonymous):

\[\cot(t)+\frac{\sin(t)}{1+\cos(t)}=\csc(t)\] I'm assuming that's what you meant? You should REALLY add parenthesis to make things clearer: cot(t) + sin(t)/[1 +cos(t)] = csc(t)

OpenStudy (anonymous):

\[\frac{\cos(t)}{\sin(t)}+\frac{\sin(t)}{1+\cos(t)}=\frac{1}{\sin(t)}\]

OpenStudy (anonymous):

Multiply by sin(t)

OpenStudy (anonymous):

\[\cos(t)+\frac{\sin^2(t)}{1+\cos(t)} = 1\]

OpenStudy (anonymous):

multiply by 1+cos(t)

OpenStudy (anonymous):

\[\cos(t)[1+\cos(t)]+\sin^2(t)=1[1+\cos(t)]\]

OpenStudy (anonymous):

\[\cos(t)+\cos^2(t)+\sin^2(t)=1+\cos(t)\]

OpenStudy (anonymous):

trig identity: \[\sin^2(t)+\cos^2(t)=1\]

OpenStudy (anonymous):

Therefore: \[\cos(t)+1 = 1+\cos(t)\]

OpenStudy (anonymous):

Just a reminder, order of operations is very important - you can't just randomly throw away parenthesis an expect people to know how to read your equation

Directrix (directrix):

I thought that the problem was intended to be: (cot t + sin t)/(1+cos t) = csc t which I don't think is an identity.

OpenStudy (anonymous):

That could be the case; I changed the order of operations in the given problem to what I think it should've been, but as it's written, like you said, the equality does not hold.